How many mL of a stock solution of 9.0 A/HNO_3 wou
Answer #1

So the amount of mols of HNO3 in the tarjet solution can be calculated as the Concentration times the volume as the following formula shows:

m=[C_{HNO_{3}}]\cdot V

and those mols have to be tha same as the stock solution used so we can do the following

m=[0.450\frac{mol}{L}]\cdot 0.380L

m=0.171 mol

0.171 mol=[9\frac{mol}{L}]\cdot V

V=\frac{0.171mol}{9\frac{mol}{L}}

V=0.019L

V=0.019L\cdot \frac{1000 ml}{1L}=19ml